# Intergroup violence in bursts or fizzles

### v10

@@ 12 Overview of cases and descriptive statistics
Table SLABEL:table:long: Overview of the groups. Occurrence of collective violence (c.v.); maximum participation in violence ($N_{C}$); time lag between the first participant and maximum participation in collective violence; burst; levels of opponents-produced turmoil ($T$) and of synchrony at first violence outbreak; number of unconditional defectors ($N_{D}$). In group 8_queensday, two focal group members disappeared from view already before the first moment of violence had started; out of the remaining 12, 6 participated in collective violence.
@@ 13.1 Mean field analysis
$H=-\sum_{i,j}w_{ij}(S_{0}+\hat{S_{i}})(S_{0}+\hat{S_{j}}),$ (4)
with $\hat{S_{i}}$ and $\hat{S_{j}}\in\{-\Delta,\Delta\}$, and taking into account the row-normalization of the adjacency matrix ($\sum_{j}^{n}w_{ij}=1$).
@@ 13.2 Critical proportion of unconditional defectors
$\Delta M/T$ =±$=\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}$
$\frac{M}{\Delta}$ =±$=\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta^{2}/T}.$ (16)
@@
±$\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta/T}$ $=S_{0}-p\Delta+(1-p)\Delta\tanh{({\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}})}$
±$\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta^{2}/T}$ $=\frac{S_{0}}{\Delta}-p\pm(1-p)\tanh{({\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}})}$
@@
$-\frac{\operatorname{arcosh}(\sqrt{\Delta^{2}(1-p)/T})}{\Delta^{2}/T}+(1-p)\sqrt{1-\frac{1}{\Delta^{2}(1-p)/T}}=\frac{S_{0}}{\Delta}-p.$ (18)

### v12

@@ 12 Overview of cases and descriptive statistics
Table SLABEL:table:long: Overview of the groups. Occurrence of collective violence (c.v.); maximum participation in violence ($N_{C}$); time lag between the first participant and maximum participation in collective violence; burst; levels of opponents-produced turmoil ($T$) and of synchrony at the first outbreak of c.v.; and, number of unconditional defectors ($N_{D}$). In case 6, two focal group members disappeared from view before the first moment of violence had started; out of the remaining 12, 6 participated in collective violence. In cases 19, 27 and 36, $NC+ND\neq n$ due to participants’ role switching. In cases 19 and 27, one of the group members defected first (deescalated; was not present at the start) but then switched to fighting. In case 36, the single fighter fell to the ground.
@@ 13.1 Mean field analysis
[…]
@@ 13.2 Critical proportion of unconditional defectors
$\Delta M/T$ $=\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}$
$\frac{M}{\Delta}$ $=\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta^{2}/T}.$ (16)
@@
$\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta/T}$ $=S_{0}-p\Delta+(1-p)\Delta\tanh{({\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}})}$
$\frac{\pm\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}}{\Delta^{2}/T}$ $=\frac{S_{0}}{\Delta}-p\pm(1-p)\tanh{({\operatorname{arcosh}{(\sqrt{\Delta^{2}(1-p)/T})}})}$
@@
$\frac{\operatorname{arcosh}(\sqrt{\Delta^{2}(1-p)/T})}{\Delta^{2}/T}+(1-p)\sqrt{1-\frac{1}{\Delta^{2}(1-p)/T}}=\frac{S_{0}}{\Delta}-p.$ (18)
@@ 13.3 Solving the Ising model computationally
Data availability
All coded video data, the R script used to produce plots from the data (Supplementary information), and a Fortran script for simulations of the Ising model are available at https://osf.io/f25nq/